# Modeling Neural Activity with Conditionally Linear Dynamical Systems

### v1

@@ 3.1 Setup
For a given trajectory $\{{\mathbf{y}}_{1:T},{\bm{u}}_{1:T}\}$, we denote as data reconstruction the mean emission $\mathbb{E}\left[{\mathbf{y}}_{1:T}\mid\hat{{\mathbf{x}}}_{1:T},{\mathbf{F}}({\bm{u}}_{1:T})\right]=({\mathbf{C}}({\bm{u}}_{1})\hat{{\mathbf{x}}}_{1},\dots,{\mathbf{C}}({\bm{u}}_{T})\hat{{\mathbf{x}}}_{T})$ from a the posterior mode $\hat{{\mathbf{x}}}_{1:T}$, computed with Kalman smoothing, given the observations ${\mathbf{y}}_{1:T}$ and parameters ${\mathbf{F}}({\bm{u}}_{1:T})$. As our primary metric, we use co-smoothing (Pei et al., 2021) to evaluate the ability of models to predict held-out single-neuron activity. Specifically, for the top 5 neurons with highest variance from the test set, we compute the coefficient of determination $R^{2}$ between the true and reconstructed single-neuron firing rate, obtained by performing data reconstruction using only the other neurons.
@@
The latent dynamical system (1a) depends on the condition ${\bm{u}}_{t}$, which can make visualizations challenging. Building on the idea that CLDS models decompose a nonlinear dynamical system into linearizations governed by ${\bm{u}}$ (see Appendix §A.2), we aim to approximate the general nonlinear system by marginalizing out ${\bm{u}}_{t}$, conditioned on ${\mathbf{x}}_{t}$. That is,
${\mathbf{x}}_{t+1}$ $=\bm{g}({\mathbf{x}}_{t})+\epsilon_{t}$
$\coloneqq\mathbb{E}_{p({\bm{u}}\mid{\mathbf{x}}_{t})}\left[{\mathbf{A}}({\bm{u}}){\mathbf{x}}_{t}+{\mathbf{b}}({\bm{u}})\right]+\epsilon_{t},$ (10)
which we define as the composite dynamical system. Intuitively, we expect this to provide a good approximation to the underlying nonlinear dynamics when ${\bm{u}}_{t}$ and ${\mathbf{x}}_{t}$ tightly co-determine each other—i.e., when the encoding $p({\mathbf{x}}_{t}\mid{\bm{u}}_{t})$ and decoding $p({\mathbf{u}}_{t}\mid{\mathbf{x}}_{t})$ conditional distributions have low variance (see Appendix §A.2.2). In practice, we estimate the expectations in (10) by computing the empirical average over ${\bm{u}}$ per binned ${\mathbf{x}}_{n}$, obtained by pooling the ${\bm{u}}_{n}$ associated with the posterior mode $\hat{{\mathbf{x}}}_{n}$ given a trajectory $\{{\mathbf{y}}_{1:T},{\bm{u}}_{1:T}\}$.
@@
For model comparison, we use as baselines (1) a standard LDS model with additive inputs of the form ${\bm{B}}{\bm{u}}_{t}$ in the latent dynamics, and (2) the LFADS (Pandarinath et al., 2018) model with controller inferred-inputs, with Gaussian observation model to fit directly to the firing rates. See Appendix §B.1 for implementation details.
@@ 3.2 Synthetic Head-Direction Ring Attractor
We start by considering a synthetic experiment of head direction neural dynamics. We conceptualize latent dynamics that capture the head direction (HD) of the animal, with attractor dynamics about a HD-dependent fixed point—see schematic in Fig. 2a. This synthetic experiment is designed to represent a nonlinear system decomposed as linear systems, per HD serving as the condition. We plot in Fig. 2b what the resulting, “composite dynamics” (see §3.1), nonlinear flow-field would be, assuming the latent state encodes the head direction exactly. The generative dynamics are an instance of a CLDS model by construction, so this synthetic example allows us to explore recovery performance.
@@
${\bm{e}}_{1}(\theta)=\begin{bmatrix}\cos(\theta)\\ \sin(\theta)\end{bmatrix},\quad{\bm{e}}_{2}(\theta)=\begin{bmatrix}-\sin(\theta)\\ \cos(\theta)\end{bmatrix},$ (11)
that describe the position on the ring and the tangent vector respectively. We design (i.e. impose) that the system converges to a stable fixed point at ${\bm{e}}_{1}(\theta)$, and at head direction $\theta$ we approximately integrate speed input along the subspace spanned by ${\bm{e}}_{2}(\theta)$. To do this, we define ${\bm{A}}(\theta)$ to be a rank-one matrix that defines a leaky line attractor, with attracting (i.e. contracting) dynamics along the orthogonal ${\bm{e}}_{1}(\theta)$. For a hyperparameter $10<\epsilon<$ 1 define:
@@
Completing the model description, we assume that the firing rate of individual neurons is given by a linear readout. For neuron $i$, the firing rate at time $t$ is:
${\mathbf{y}}_{t,i}={\bm{C}}_{i,:}(\theta_{t})^{\top}{\mathbf{x}}_{t}+\bm{\omega}_{t}$ (13)
@@
where ${\bm{C}}_{i,:}$ is sinusoidal bump tuning curve function (see App. §B.2). Note that we set ${\mathbf{d}}({\bm{u}}_{t})=\bm{0}$. Finally, we sampled trials of length $T=100$ with $M=10$ neurons, generating the evolution of the heading direction as a random walk, $\theta_{t}\sim\mathcal{N}(\theta_{t-1},0.5^{2})$, and initialize at the origin ${\mathbf{x}}_{1}\sim\mathcal{N}(0,1)$.
@@
We report our recovery results in Fig. 2, fixing the decoding matrix ${\mathbf{C}}(\cdot)$ to a known value as to avoid non-identifiability considerations. We refer to Appendix §B.2 for recovery plots of ${\mathbf{C}}$ when fitted. First, we observed that we can generally recover the nonlinear flow-field, plotting in Fig. 2c the composite dynamics obtained from the posterior trajectories. This paints an activity-based portrait of the dynamics, and our ability to accurately estimate the flow-field around a given point under this method depends on how many posterior samples pass by it. We thus indicate, by gray-scale shading the flow-field arrows, the fraction of posterior samples pooled per bin as a fraction of the highest bin count. Second, for a more parameter-based account of the recovery, we plot in Fig. 2d-e the varying biases ${\mathbf{b}}(\theta)$ and dynamics matrices ${\mathbf{A}}(\theta)$ as functions of the head direction $\theta$—we recovered with high-fidelity the true parameters. This recovery translated into the properties of the dynamics such as the recovered eigenvalues of ${\mathbf{A}}(\theta)$ in Fig. 2f. Finally, we observed that the test data single-neuron reconstruction (Fig. 2g) recovers the true observations, and the model was able to accurately ($R^{2}=0.86$) reconstruct a held out neuron from this test-set through co-smoothing.
3.3 Mouse Head-Direction Circuit Dynamics
@@
Next, we turned to the analysis of antero-dorsal thalamic nucleus (ADn) recordings from Peyrache et al. (2015) of the mouse HD system in mice foraging in an open environment (Fig. 3a). We considered neural activity from the “wake” period, binned in 50ms time-bins, then processed to firing rates and separated into 10s trials. As with the synthetic experiment of the previous section, we treat the recorded head-direction $\theta_{t}\in[0,2\pi)$ as conditions ${\bm{u}}_{t}=\theta_{t}$.
@@
Finally, we analyzed the learned latent dynamics. Like the synthetic example, we identified a ring attractor structure (Fig. 3d). Now, unlike the synthetic example, we observed that this ring attractor is composed of HD-dependent fixed points as opposed to line attractors, as per the eigenvalues of ${\mathbf{A}}(\theta)$ in Fig. 3e.
@@ 3.4 Macaque Center-Out Reaching Task
Finally, we analyzed neural recordings of dorsal premotor cortex (PMd) in macaques performing center-out reaching task (Fig. 4a) from N. Even-Chen, B. Sheffer et al. (2019). In contrast to the previous experimental conditions, we consider here two-dimensional conditions $\bm{{\bm{u}}}^{k}_{t}=(\theta^{k},z_{t})$, where $\theta^{k}\in[0,2\pi)$ is the instructed reach angle, constant per trial $k$, and $z_{t}\in\{0,1\}$ indicates the task reach condition (see Fig. 4b) set at 00 during the delay and $1$ at $100$ms past the go-cue, at the onset of the movement-related firing rate ramp Fig. 4a-(right). Discrete-valued conditions, such as the reach onset $z_{t}\in\{0,1\}$, are considered as supported on a continuous interval. The correlation between such discrete points is determined by the length-scale parameter $\kappa$, which we’ve set to $\kappa=0.5$ from a hyperparameter search. More details on hyperparameters and data-preprocessing can be found in Appendix §B. Finally, we use a fixed emission matrix ${\bm{C}}$ and let the latent dynamics capture the dependency on experimental conditions through ${\mathbf{A}}({\bm{u}})$ and ${\mathbf{b}}({\bm{u}})$.
We found the latent dynamics to encode the conditions through attracting fixed-points during both the delay and reach periods. We show in Fig. 4c the projection of the $D=5$ latent dimensions along the $3$-dimensional subspace most aligned (i.e. best decoding) with the experimental conditions, following similar analyses from (N. Even-Chen, B. Sheffer et al., 2019)—we observed clear aligned rings of fixed points from delay to reach. In CLDS models, we obtain the fixed points by simply solving for ${\bm{x}}^{*}({\bm{u}})$ satisfying $({\bm{I}}-{\mathbf{A}}({\bm{u}})){\bm{x}}^{*}={\mathbf{b}}({\bm{u}})$ for any ${\bm{u}}$, in contrast to numerical fixed-point methods usually employed (Sussillo & Barak, 2013).
We performed co-smoothing (see §3.1) to evaluate the model. We recovered with good accuracy single held-out neurons from the validation set excluded from training (Fig. 4d). We then compared the performance of the CLDS against the LDS and LFADS models, exploring further how each fares in low-data regimes. We report in Fig. 4e the co-smoothing $R^{2}$ per model, computed as a function of the number of trials used in each reach-angle $\theta^{k}$, averaged over 5 random seeds. We found that the CLDS outperformed both models consistently, with the highest difference at $1$ training trial per condition. While the LFADS model showed progressively better performance that did not plateau yet, it nonetheless underperformed in these low data regimes.
@@
Figure 4: Macaque reaching experiment. (a) Task schematic (left) and population-averaged firing rates per trial (right). (b) 2D conditions, with trial orientation $\theta$, and reach variable $z_{t}\in\{0,1\}$ switching at ramp onset. (c) 3D projection of the 5 dimensional latents used, projected as to align best with condition decoding. We show the model fixed-points per reach angle $\theta$ and reach condition $z$, plotted over posterior mean trajectories per trial. (d) Co-smoothing reconstruction of single held-out neurons from the test-set. (e) Co-smoothing $R^{2}$ per model as a function of the number of trials used per reach angle during training. Error bars indicated std. around the mean over 5 random initialization seeds.
@@ 4 Related work
CLDS models capture the dependence of neural responses ${\mathbf{y}}_{1:T}$ on continuous experimental conditions ${\bm{u}}_{1:T}$. Nejatbakhsh et al. (2023) investigated a very similar problem, focusing on single-trial responses ${\mathbf{y}}_{k}$ to continuous experimental conditions ${\bm{u}}_{k}\in\mathcal{U}$. They use a conditional Gaussian model for responses ${\mathbf{y}}_{k}$ given conditions ${\bm{u}}_{k}$
${\mathbf{y}}_{k}\mid{\bm{u}}_{k}\sim\mathcal{N}({\mathbf{y}};\mu({\bm{u}}_{k}),\Sigma({\bm{u}}_{k}))$ (15)
and they place Gaussian process and Wishart process (Wilson & Ghahramani, 2011) priors on the mean and covariance functions. Concretely, they posit that
$\mu(\cdot)$ $\sim\mathcal{GP}^{M}(0,k_{\mu}),\quad$ (16a)
$\Sigma({\bm{u}})$ $={\mathbf{U}}({\bm{u}}){\mathbf{U}}({\bm{u}})^{\top}+\Lambda({\bm{u}})$ (16b)
@@
For a single time step $T=1$, $t=1$, our system in (1) reads
$p({\mathbf{x}}_{1}\mid{\bm{u}}_{1})$ $=\mathcal{N}({\mathbf{x}}_{1};{\mathbf{m}}({\bm{u}}_{1}),Q_{1})$
$p({\mathbf{y}}_{1}\mid{\mathbf{x}}_{1},{\bm{u}}_{1})$ $=\mathcal{N}\left({\mathbf{y}}_{1};{\mathbf{C}}({\bm{u}}_{1}){\mathbf{x}}_{1}+{\mathbf{d}}({\bm{u}}_{1}),R\right),$
Assuming a degenerate prior that ${\mathbf{m}}({\bm{u}}_{1})=\bm{0}$, the marginal distribution of ${\mathbf{y}}_{1}$ conditioned on ${\bm{u}}_{1}$ equals
@@
which can be compared with (15) and (16). We observe that the models are essentially equivalent with $\mu({\bm{u}})={\mathbf{d}}({\bm{u}})$, and with the CLDS emission matrix ${\mathbf{C}}({\bm{u}})$ serving as the Wishart process prior decomposition matrix ${\mathbf{U}}({\bm{u}})$, right-scaled by $Q_{1}^{1/2}\in\mathbb{R}^{D\times D}$. This makes the parameter $p=D$ now bear meaning as the dimensionality of the latents ${\mathbf{x}}\in\mathbb{R}^{D}$. Thus, we can view CLDS models as a direct extension of Wishart process models that capture condition-dependent dynamics across multiple time steps.
Markovian GPs Latent GP models (Lawrence, 2007; Wang et al., 2005), such as the foundational model of GPFA by Yu et al. (2009), are widely used in neuroscience. Here we show that one can view MAP inference in a CLDS model as optimizing a kernel that defines a latent GP prior. While GPFA is not a dynamical system model, Yu et al. (2009), as well as Turner & Sahani (2007), detail how the the stationary dynamics of an AR-1 process (i.e. linear dynamical system) can be expressed as draws from a GP. More generally, all stationary, real-valued, and finitely differentiable GPs admit a representation in terms of linear state space models (Dowling et al., 2021, 2023). The main departure with our work is that we additionally place a GP prior on the parameters (coefficients) of an LDS, allowing the dynamics to vary across conditions and across time. For a fixed set of LDS parameters, ${\bm{F}}$, and experimental covariates, ${\bm{u}}_{1:T}$, the distribution of latent states in a CLDS are jointly Gaussian. Thus, a set of LDS parameters induces a (generally non-stationary) GP prior on the latent trajectories. In this view, the GP prior we place over the parameters of the LDS can be seen as a hyperprior over the latent dynamical process.
@@
A second class of relevant models generalizing the LDS are Switching LDS models (SLDS; Murphy 1998; Pavlovic et al. 2000; Petreska et al. 2011). SLDS models consist of a discrete latent state ${\textnormal{z}}_{t}$ with Markov chain dynamics dictating the dynamics matrix ${\bm{A}}^{({\textnormal{z}}_{t})}$. This switching behavior can be mimicked in our setting if the condition space is discrete (see, e.g., §3.4). We can take the relationship a step further by embedding the discrete process in the continuous dynamics parameter space of ${\bm{A}}^{({\textnormal{z}}_{t})}\in\mathbb{R}^{D\times D}$. Under this lens and in a similar line of thinking as with Markovian GPs, we show in Appendix §A.3 how a one-dimensional SLDS model with latent dynamics
$p({\textnormal{z}}_{t+1}=i\mid{\textnormal{z}}_{t}=j)$ $=P_{ij},\quad{\textnormal{x}}_{t+1}=a^{({\textnormal{z}}_{t})}{\textnormal{x}}_{t}+\epsilon_{t}$
@@
${\textnormal{a}}(\cdot)\sim\mathcal{GP}\left(\pi^{\top}\bm{a},\bm{a}^{\top}\left(P^{\left|t_{j}-t_{i}\right|}\mathrm{diag}(\pi)-\pi\pi^{\top}\right)\bm{a}\right),$
over time conditions ${\bm{u}}_{t}=t$, for $\bm{a}$ the vector of values taken by $a^{(z_{i})}$ and $\pi$ the stationary distribution of the ${\textnormal{z}}_{t}$ discrete state process. Finally, in a similar vein, Geadah et al. (2024) consider the discrete states ${\textnormal{z}}_{t}$ dictating the dynamics to live on a continuous support. However, they do not leverage this continuity in the parameters ${\mathbf{A}}(\cdot)$ themselves.
The recurrent SLDS (rSLDS) model (Linderman et al., 2017) takes an important departure from the SLDS by leveraging the continuous latent states ${\mathbf{x}}_{t}$ to guide the discrete state transitions. Smith et al. (2021) use this dependency but turn to the linearization of nonlinear systems, using ${\mathbf{x}}$-space fixed points as guide for the linear dynamics. In contrast, we linearize based on observed external conditions.

### v2

@@ 3.1 Setup
For a given trajectory $\{{\mathbf{y}}_{1:T},{\bm{u}}_{1:T}\}$, we denote as data reconstruction the mean emission $\mathbb{E}\left[{\mathbf{y}}_{1:T}\mid\hat{{\mathbf{x}}}_{1:T},{\mathbf{F}}({\bm{u}}_{1:T})\right]=({\mathbf{C}}({\bm{u}}_{1})\hat{{\mathbf{x}}}_{1},\dots,{\mathbf{C}}({\bm{u}}_{T})\hat{{\mathbf{x}}}_{T})$ from a the posterior mode $\hat{{\mathbf{x}}}_{1:T}$, computed with Kalman smoothing, given the observations ${\mathbf{y}}_{1:T}$ and parameters ${\mathbf{F}}({\bm{u}}_{1:T})$. As our primary metric, we use co-smoothing (PeiYe2021NeuralLatents, ) to evaluate the ability of models to predict held-out single-neuron activity. Specifically, for the top 5 neurons with highest variance from the test set, we compute the coefficient of determination $R^{2}$ between the true and reconstructed single-neuron firing rate, obtained by performing data reconstruction using only the other neurons.
@@
The latent dynamical system (1a) depends on the condition ${\bm{u}}_{t}$, which can make visualizations challenging. Building on the idea that CLDS models decompose a nonlinear dynamical system into linearizations governed by $\bm{u}$} (see App. §A.2), we aim to approximate the general nonlinear system by marginalizing out ${\bm{u}}_{t}$, conditioned on ${\mathbf{x}}_{t}$. That is,
${\mathbf{x}}_{t+1}$ $=\bm{g}({\mathbf{x}}_{t})+\epsilon_{t}\coloneqq\mathbb{E}_{p({\bm{u}}\mid{\mathbf{x}}_{t})}\left[{\mathbf{A}}({\bm{u}}){\mathbf{x}}_{t}+{\mathbf{b}}({\bm{u}})\right]+\epsilon_{t},$ (7)
which we define as the composite dynamical system. Intuitively, we expect this to provide a good approximation to the underlying nonlinear dynamics when ${\bm{u}}_{t}$ and ${\mathbf{x}}_{t}$ tightly co-determine each other—i.e., when the encoding $p({\mathbf{x}}_{t}\mid{\bm{u}}_{t})$ and decoding $p({\mathbf{u}}_{t}\mid{\mathbf{x}}_{t})$ conditional distributions have low variance (see App. §A.2.2). In practice, we estimate the expectation in (7) by averaging over the ${\bm{u}}_{n}$ associated with binned values $\hat{{\mathbf{x}}}_{n}$ of the posterior mode given a trajectory $\{{\mathbf{y}}_{1:T},{\bm{u}}_{1:T}\}$. Thus our ability to accurately estimate the flow-field around a given point under this method depends on how many posterior samples pass by it, which we report alongside the composite dynamics plots.
@@
For model comparison, we use as baselines (1) a standard LDS model and (2) a more flexible gpSLDS (hu2024modeling, ) (implemented in a way to include linear boundaries to encompass the rSLDS linderman:rSLDS ), both with additive inputs of the form ${\bm{B}}{\bm{u}}_{t}$ in the latent dynamics, and (3) LFADS (LFADS, ) model with controller inferred-inputs as a fully nonlinear alternative. For all, we consider Gaussian observation model to fit directly to the firing rates. See Appendix §B.1 for implementation details.
@@ 3.2 Synthetic head-direction ring attractor
We start by considering a synthetic experiment of head direction neural dynamics. We conceptualize latent dynamics that capture the head direction (HD) of the animal, with attractor dynamics about a HD-dependent fixed point—see schematic in Fig. 2a. This synthetic experiment is designed to represent a nonlinear system decomposed as linear systems, per HD serving as the condition. We plot in Fig. 2b what the resulting, “composite dynamics” (see §3.1), nonlinear flow-field would be, assuming the latent state encodes the head direction exactly. The generative dynamics are an instance of a CLDS model by construction, so this example allows us to explore recovery performance.
@@
${\bm{e}}_{1}(\theta)=\begin{bmatrix}\cos(\theta)&\sin(\theta)\end{bmatrix},\quad{\bm{e}}_{2}(\theta)=\begin{bmatrix}-\sin(\theta)&\cos(\theta)\end{bmatrix}$, (8)
that describe the position on the ring and the tangent vector respectively. We design (i.e. impose) that the system converges to a stable fixed point at ${\bm{e}}_{1}(\theta)$, and at head direction $\theta$ we approximately integrate speed input along the subspace spanned by ${\bm{e}}_{2}(\theta)$. To do this, we define ${\bm{A}}(\theta)$ to be a rank-one matrix that defines a leaky line attractor, with attracting (i.e. contracting) dynamics along the orthogonal ${\bm{e}}_{1}(\theta)$. For a hyperparameter $0<\epsilon<1$ define:
@@
Completing the model description, we assume that the firing rate of individual neurons is given by a linear readout as ${\mathbf{y}}_{t,i}={\bm{C}}_{i,:}(\theta_{t})^{\top}{\mathbf{x}}_{t}+\bm{\omega}_{t}$ for neuron $i$ at time $t$, with ${\mathbf{d}}({\bm{u}}_{t})=\bm{0}$ and ${\bm{C}}_{i,:}$ is sinusoidal bump tuning curve function (see App. §B.2). Finally, we sampled trials of length $T=100$ with $N=10$ neurons, generating the evolution of the heading direction as a random walk, $\theta_{t}\sim\mathcal{N}(\theta_{t-1},0.5^{2})$, and initialize at the origin ${\mathbf{x}}_{1}\sim\mathcal{N}(0,1)$.
@@
We report our recovery results in Fig. 2, fixing the decoding matrix ${\mathbf{C}}(\cdot)$ to a known value as to avoid non-identifiability considerations. We refer to App. §B.2 for recovery plots of $\mathbf{C}$} when fitted. First, we observed that we can generally recover the nonlinear flow-field, plotting in Fig. 2c the composite dynamics obtained from the posterior trajectories. Second, for a more parameter-based account of the recovery, we plot in Fig. 2d-e the varying biases ${\mathbf{b}}(\theta)$ and dynamics matrices ${\mathbf{A}}(\theta)$ as functions of the head direction $\theta$—we recovered with high-fidelity the true parameters. This recovery translated into the properties of the dynamics such as the recovered eigenvalues of ${\mathbf{A}}(\theta)$ in Fig. 2f. Third, we observed that the test data single-neuron reconstruction (Fig. 2g) recovers the true observations, and the model was able to accurately ($R^{2}=0.86$) reconstruct a held out neuron from this test-set through co-smoothing. Finally, we report in Appendix §B.2 a study on the impact of observational noise on our quality of fit—we find reliable recovery even as the signal-to-noise ratio degrades.
3.3 Neural circuit model of ring-attractor dynamics
We provide an alternate synthetic ring attractor experiment based on the neuroscience literature Hulse2020 . It is meant to test the CLDS and its inference in a synthetic data modality with known underlying computation but which, importantly, is not generated from a CLDS model.
We write a model of continuous ring attractor dynamics with bumps of activity integrating angular velocity Zhang1996 . In this model, high-dimensional and non-normal attractor dynamics are implemented through HD-preferential units arranged in a “ring”, with short-range excitation and long-range inhibition forming a bump of activity for a specific unit in a specific HD, and with a HD velocity $v(t)=\dot{\theta}(t)$ integrator causing a shift in the location of the bump. Let ${\textnormal{y}}(\phi,t)$ denote the network activity of cells with preferred head-direction $\phi\in[0,2\pi)$ at time $t$. The dynamics follow the stochastic partial differential equation
$\frac{\partial}{\partial t}{\textnormal{y}}=\tau\left[-{\textnormal{y}}+f(w*{\textnormal{y}})+v\,\frac{\partial}{\partial\phi}{\textnormal{y}}\right]+\sigma\xi$ (10)
for $\xi(\phi,t)$ a white noise process on the ring. Here, $f$ is a nonlinearity, $w$ a “Mexican-hat” convolution filter, and $\tau>0$ a time constant. For implementation purposes, y is discretized to ${\mathbf{y}}_{t}\in\mathbb{R}^{N}$, $N=32$, for $t\in\{1,\dots,T\}$, with each unit ${\mathbf{y}}]_{i}$ having preferred directions $\phi_{i}$ regularly spanning the interval $0,2\pi)$. See model and implementation details in Appendix §B.3. We use $\mathbf{y}$} as our observations and head-direction $\theta$ as conditions again, generated the same way as in our previous synthetic HD model.
The underlying model captures high-dimensional ring attractor dynamics. When fitting a CLDS model, we expect the low-dimensional latent dynamics to capture the ring-attractor structure and the nonlinearity of the dynamics, with a linear emission model bringing it back to the high-dimensional activity. This is precisely what we found when fitting the CLDS—see Appendix §B.3. The results make us confident that the CLDS can capture this nonlinear structure even under model mismatch.
3.4 Mouse head-direction dynamics
@@
Next, we turned to the analysis of antero-dorsal thalamic nucleus (ADn) recordings from Ref. Peyrache2015 of the mouse HD system in mice foraging in an open environment (Fig. 3a). We considered neural activity from the “wake” period, binned in 50ms time-bins, then processed to firing rates and separated into 10s trials. As with the synthetic experiment of the previous section, we treat the recorded head-direction $\theta_{t}\in[0,2\pi)$ as conditions ${\bm{u}}_{t}=\theta_{t}$.
@@
Finally, we analyzed the learned latent dynamics. Like the synthetic example, we identified a ring attractor structure (Fig. 3d). Unlike the synthetic example, per the eigenvalues of ${\mathbf{A}}(\theta)$ in Fig. 3e, this ring attractor is composed of HD-dependent fixed points as opposed to line attractors.
@@ 3.5 Macaque center-out reaching task
Figure 4: Macaque reaching experiment. (a) Task schematic (left) and population-averaged firing rates per trial (right). (b) 2D conditions, with trial orientation $\theta$, and reach variable $z_{t}\in\{0,1\}$ switching at ramp onset. (c) 3D projection of the 5 dimensional latents used, projected as to align best with condition decoding. We show the model fixed-points per reach angle $\theta$ and reach condition $z$, plotted over posterior mean trajectories per trial. (d) Co-smoothing reconstruction of single held-out neurons from the test-set. (e) Co-smoothing $R^{2}$ per model as a function of the number of trials used per reach angle during training. Error bars indicated std. around the mean over 5 initialization seeds.
Finally, we analyzed neural recordings of dorsal premotor cortex (PMd) in macaques performing center-out reaching task (Fig. 4a) from Ref. EvenChen2019 . In contrast to the previous experimental conditions, we consider here two-dimensional conditions $\bm{{\bm{u}}}^{k}_{t}=(\theta^{k},z_{t})$, where $\theta^{k}\in[0,2\pi)$ is the instructed reach angle, constant per trial $k$, and $z_{t}\in\{0,1\}$ indicates the task reach condition (see Fig. 4b) set at 0 during the delay and $1$ at 100100ms past the go-cue, at the onset of the movement-related firing rate ramp Fig. 4a-(right). Discrete-valued conditions, such as the reach onset $z_{t}\in\{0,1\}$, are considered as supported on a continuous interval. The correlation between such discrete points is determined by the length-scale parameter $\kappa$, which we’ve set to $\kappa=0.5$ from a hyperparameter search. More details on hyperparameters and data-preprocessing can be found in Appendix §B. Finally, we use a fixed emission matrix $\bm{C}$} and let the latent dynamics capture the dependency on experimental conditions through ${\mathbf{A}}({\bm{u}})$ and ${\mathbf{b}}({\bm{u}})$.
We found the latent dynamics to encode the conditions through attracting fixed-points during both the delay and reach periods. We show in Fig. 4c the projection of the $D=5$ latent dimensions along the $3$-dimensional subspace most aligned (i.e. best decoding) with the experimental conditions, following common analyses EvenChen2019 —we observed clear aligned rings of fixed points from delay to reach. In CLDS models, we obtain the fixed points by simply solving for ${\bm{x}}^{*}({\bm{u}})$ satisfying $({\bm{I}}-{\mathbf{A}}({\bm{u}})){\bm{x}}^{*}={\mathbf{b}}({\bm{u}})$ for any $\bm{u}$}, in contrast to numerical fixed-point methods usually employed (sussillo2013opening, ).
We performed co-smoothing (see §3.1) to evaluate the model. We recovered with good accuracy single held-out neurons from the validation set excluded from training (Fig. 4d). We then compared the performance of the CLDS against the baseline models, exploring further how each fares in low-data regimes. We report in Fig. 4e the co-smoothing $R^{2}$ per model, computed as a function of the number of trials used in each reach-angle $\theta^{k}$, averaged over 5 random seeds. We found that the CLDS outperformed other models, displaying the highest difference at $1$ training trial per condition. Introducing more training trials per condition, the gpSLDS model sometimes approached the performance of the CLDS. While the LFADS model showed progressively better performance that did not plateau yet, it nonetheless underperformed in these low data regimes.
@@ 4 Related work
CLDS models capture the dependence of neural responses ${\mathbf{y}}_{1:T}$ on continuous experimental conditions ${\bm{u}}_{1:T}$. Ref. Nejatbakhsh:2023 investigated a very similar problem, focusing on single-trial responses ${\mathbf{y}}_{k}$ to continuous experimental conditions ${\bm{u}}_{k}\in\mathcal{U}$. They use a conditional Gaussian model for responses ${\mathbf{y}}_{k}$ given conditions ${\bm{u}}_{k}$
${\mathbf{y}}_{k}\mid{\bm{u}}_{k}\sim\mathcal{N}({\mathbf{y}};\mu({\bm{u}}_{k}),\Sigma({\bm{u}}_{k}))$ (12a)
and they place Gaussian process and Wishart process wilson2011generalised priors on the mean and covariance functions
$\mu(\cdot)$ $\sim\mathcal{GP}^{M}(0,k_{\mu}),\qquad\Sigma({\bm{u}})={\mathbf{U}}({\bm{u}}){\mathbf{U}}({\bm{u}})^{\top}+\Lambda({\bm{u}})$ (12b)
@@
For a single time step $t=T=1$ and assuming a degenerate prior ${\mathbf{m}}({\bm{u}}_{1})=\bm{0}$, the marginal distribution of ${\mathbf{y}}_{1}$ conditioned on ${\bm{u}}_{1}$ in our system (1) reads
@@
which can be compared with (12). The models are analogous with $\mu({\bm{u}})={\mathbf{d}}({\bm{u}})$ and with the CLDS emission matrix ${\mathbf{C}}({\bm{u}})$ serving as the Wishart process prior decomposition matrix ${\mathbf{U}}({\bm{u}})$, right-scaled by $Q_{1}^{1/2}\in\mathbb{R}^{D\times D}$. This makes the parameter $p=D$ now bear meaning as the dimensionality of the latents ${\mathbf{x}}\in\mathbb{R}^{D}$, akin to factor analysis. Thus, we can view CLDS models as a direct extension of Wishart process models that capture condition-dependent dynamics across multiple time steps.
Markovian GPs Latent GP models GPLVM ; NIPS2005_ccd45007 , such as GPFA Yu2009 , are widely used in neuroscience. One can view MAP inference in a CLDS model as optimizing a kernel that defines a latent GP prior. Taking this a step further, the stationary dynamics of an AR-1 process (i.e. linear dynamical system) can be expressed as draws from a GP Yu2009 ; Turner2007 . Vice versa, all stationary, real-valued, and finitely differentiable GPs admit a representation in terms of linear state space models (dowling2021hida, ; pmlr-v202-dowling23a, ), a relationship that has proven useful for both modeling and aiding GP inference (pmlr-v202-dowling23a, ; Hartikainen2010, ; Li:2024, ). We expand upon this direction by allowing the dynamics to vary not only across time but over conditions too. For a fixed set of LDS parameters at experimental covariates, ${\bm{F}}({\bm{u}}_{1:T})$, the distribution of latent states in a CLDS are jointly Gaussian. Thus, a set of LDS parameters induces a (generally non-stationary) GP prior on the latent trajectories. In this view, the GP prior we place over the parameters of the LDS can be seen as a hyperprior over the latent dynamical process.
@@
A second class of relevant models generalizing the LDS are Switching LDS models (SLDS; murphy1998learning ; NIPS2000_ca460332 ; Petreska2011 ). SLDS models consist of a discrete latent state ${\textnormal{z}}_{t}$, with finite Markov chain dynamics, dictating the dynamics matrix ${\bm{A}}^{({\textnormal{z}}_{t})}$. This switching behavior can be mimicked in our setting if the condition space is discrete (see, e.g., §3.5). We can take the relationship a step further by embedding the discrete process in the continuous dynamics parameter space of ${\bm{A}}^{({\textnormal{z}}_{t})}\in\mathbb{R}^{D\times D}$. In a similar line of thinking as with Markovian GPs, we show in Appendix §A.3 how a one-dimensional SLDS model of dynamics
$p({\textnormal{z}}_{t+1}=i\mid{\textnormal{z}}_{t}=j)$ $=P_{ij},\quad{\textnormal{x}}_{t+1}=a^{({\textnormal{z}}_{t})}{\textnormal{x}}_{t}+\epsilon_{t},$
@@
${\textnormal{a}}(\cdot)\sim\textstyle\mathcal{GP}\left(\pi^{\top}\bm{a},\bm{a}^{\top}\left(P^{\left|t_{j}-t_{i}\right|}\mathrm{diag}(\pi)-\pi\pi^{\top}\right)\bm{a}\right),$
over time conditions ${\bm{u}}_{t}=t$, for $\bm{a}$ the vector of values taken by $a^{(z_{i})}$ and $\pi$ the stationary distribution of the ${\textnormal{z}}_{t}$ discrete state process. In a similar vein, previous work has considered the discrete states ${\textnormal{z}}_{t}$ dictating the dynamics to live on a continuous support geadah2024parsing , however without leveraging this continuity in the parameters ${\mathbf{A}}(\cdot)$ themselves.
The recurrent SLDS (rSLDS) model (linderman:rSLDS, ; zoltowski:2020a:rSLDS, ; brenner2024almostlinear, ) takes an important departure from the SLDS by leveraging the continuous latent states ${\mathbf{x}}_{t}$ to guide the discrete state transitions. Recent work JSLDS use this dependency but turn to the linearization of nonlinear systems with $\mathbf{x}$}-space fixed points as guide for the linear dynamics. In contrast, we linearize based on observed conditions (see App. A.2).